OpenStudy (anonymous):

Hello Need help. stats problem. a coin is flliped until head appear. what is the probability that a head appears on an odd-numbred flip? Please help. I would really appreciate it. thanks.

OpenStudy (anonymous):

1/2 + 1/2*1/2*1/2 + (1/2)^5.... (events appears on 1st =1/2 , appears on 3rd = 1/2*1/2*1/2, on 5th = (1/2)^5 ..so on) 1/2 * 1/(1-1/4) = 2/3